the sum of a positive and negative number is 0 _________.

the quotient formed by any nonzero number divided by 0 is __________, and the quotient formed by 0 divided by any nonzero is __________.
i have to show an example of the two questions also.

Answer by RAY100(1637)

sum of positive and negative number,,,,"is difference of the numbers with sign of greater number",,,,,ex (3) + (-2) = +1,,,,,difference (1),,,sign (+),,,,
example (-3) + (2) = -1,,,difference(1),,, sign (-)
.
Number divided by zero is " undefined",,,ex 2/0 = UD,,,,,-2/0=UD
.
zero divided by any number is "zero",,,,,ex.,,,0/2=0,,,0/-2=0
.